Find the EAR in each of the following cases. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16. Use 365 days in a year.)

Stated Rate (APR) Number of Times
Compounded
Effective Rate (EAR)
8.25 % Quarterly %
17.75 Monthly %
14.25 Daily %
10.25 Semiannually %
Answer #1
Stated Rate (APR) [%] Number of Times Effective Rate (EAR)
Compounded Calculation Rate [%]
8.25 Quarterly =(1+0.0825/4)^4-1 = 8.51%
17.75 Monthly =(1+0.1775/12)^12-1 = 19.27%
14.25 Daily =(1+0.1425/365)^365-1 = 15.31%
10.25 Semiannually =(1+0.1025/2)^2-1 = 10.51%